Assume technological growth, A, occurs. This may be like the tech boom of the 1990s. I have modified the price and wage setting curves to include expectations about prices and about technological change.

P= (1+)*

W = AⓇPF(u, 2)

You over hear a conversation in which one person suggests that such technological growth will result in nothing but a higher real wage, and no change to the unemployment rate. The other person suggests that there may, indeed be a decrease in unemployment. They turn to you, having just become an expert in this due to your recent Macro class, to settle the argument.

a) Using graphs and explanations show and explain how (under which conditions) each may be correct.

b) Using graphs and explanations show and explain how (under which conditions) there are alternatives

that neither have thought of. [More points for two alternatives or more].

c) There is one big assumption behind a lot of this. Recall Dilbert(comic strip). Explain/discuss.
Answer #1

Solution:

The given technological function P = (1+ µ ) W / A . Then we have A = (1+ µ ) W / P where W/P is the real wage. When A increases then real wage will also increase. If the wages rises higher than the productivity then price of the goods will increase so it will result inflation of the economy. There is a positive relation between the real wage and inflation.

Now we consider the Phillips curve. It shows the relationship between the inflation and unemployment rate. This is shown as in our figure.

Inflation LR PC A SR PC Unemployment Rate U U U

SR PC is the short run Phillips curve it is downward sloping implies that if inflation increases then unemployment decreases. LR PC is the long run Phillips curve it is vertical at certain level of employment.

In short run all the resources are not fully utilised so a technological growth will increase the real wage that will in turn increase the inflation and then level of unemployment will fall is the economy. Consider the economy is at point B now due to improvement in technology it shift to point A with reduction in unemployment from U' to U" and rise in real wage and inflation from P to P'.

But consider the situation where the resources are fully used so an improvement in the technology will lead to rise in wage and inflation but the employment is fixed at U.

We find that the recall rate of separated workers who enter unemployment is coun-tercyclical, if we focus on those workers who remain unemployed and do not leave labor force, if we include those who drop out of the labor force; hence, it appears that, in recessions, some discouraged unemployed workers leave the labor force due to a drop
in recall chances, and the share of recalls in total hires rises because other hires decline much faster than recalls.